Please help me find the dimensions of a rectangle that has an area of 10 and a diagonal of length 5. Thank you!!
Answer by venugopalramana(3286)

Please help me find the dimensions of a rectangle that has an area of 10 and a diagonal of length 5. Thank you!!
LB=10
SQRT(L^2+B^2)=5...FROM PYTHOGARUS THEOREM
L^2+B^2=25
(L+B)^2=L^2+B^2+2LB=25+20=45
L+B=SQRT(45).............................I
(L-B)^2=L^2+B^2-2LB=25-20=5
L-B=SQRT(5)...................................II
EQN.I+EQN.II...GIVES
2L=SQRT(45)+SQRT(5)
L={SQRT(45)+SQRT(5)}/2=4.472
EQN.I-EQN.II GIVES
B={SQRT(45)-SQRT(5)}/2=2.236